# VaLID: Verification as Late Integration of Detections for LiDAR-Camera Fusion

### v2

@@ IV-A Method Overview
IV-B Implementation
To implement our late-fusion method, we combine two specialized LiDAR detectors (PVRCNN and TED) with various camera models: a specialized model (MonoDETR), a general-purpose model (YOLO-NAS), and an open-vocabulary model (GroundingDINO). The experiments are conducted on the KITTI dataset. We remove dataset images for which ground truth is not available, as well as those that have already been used to train our baseline models. This leaves 3769 images available for our experiments comprising the subset that the dataset providers call validation. We split these images into training and testing subsets. For PVRCNN, these images generate a training set of 12,303 LiDAR bounding box samples and test set of 12,817 bounding boxes. For TED-S, we obtain 9,016 training and 9,406 testing bounding boxes.
The final inputs to our model include details such as LiDAR bounding box dimensions, confidence scores, and IoUs from the camera methods. These inputs pass through a two-layer multi-layer perceptron with ReLU [24] activation function and a last sigmoid layer. This output decides whether to keep each LiDAR box or discard the detection. Since our focus is on reducing false positives without significantly disturbing the true positive count, we use class weights of 1:10 for negative and positive classes to maintain high recall. The model is optimized using a binary cross-entropy loss function, which compares the sigmoid outputs against the ground truth labels to predict the acceptance or rejection of a given LiDAR detection. We train our models for 50 epochs with an Adam learning rate of 0.0001. During the inference time, we rescore the accepted LiDAR box confidence score by multiplying it by the output sigmoid prediction probability, refining the final detection confidence.
@@ V-A False Positives Reduction
Models Modality Model Type Easy Moderate Hard
PVRCNN[32] L Specialized 64.25 / 72.00 57.00 / 66.58 52.00 / 63.34
Ours (PVRCNN+YOLO-NAS) LC Late-Fusion 67.09 / 73.54 59.79 / 68.57 53.74 / 64.69
Ours (PVRCNN+GDino) LC Late-Fusion 65.64 / 72.02 58.42 / 67.30 52.82 / 63.42
Ours (PVRCNN+YOLO-NAS+GDino) LC Late-Fusion 67.28 / 73.97 60.36 / 68.01 54.31 / 63.58
TABLE III: Perfomance on the Average Precision (AP) metric for our method on all difficulty sets on KITTI-val for the Pedestrian class. Our late fusion method consistently outperforms the original specialized methods in 3DAP across all the difficulty levels. The color coding in yellow emphasizes the degree of improvement on 3DAP (primary comparison metric) over the baselines.
@@ V-B Average Precision
Late fusion with general and open-vocabulary camera models (YOLO-NAS and GroundingDINO) perform as well as the tested special-purpose camera model (MonoDETR). This is surprising since the baseline AP of these detectors is much lower than the specially trained model. This provides evidence that general-purpose camera detectors can be used to enhance special-purpose LiDAR detectors.

### v3

@@ V-A False Positives Reduction
Models Easy Moderate Hard
PVRCNN[28] 64.25 / 72.00 57.00 / 66.58 52.00 / 63.34
Ours (PV+YOLO) 67.09 / 73.54 59.79 / 68.57 53.74 / 64.69
Ours (PV+GDino) 65.64 / 72.02 58.42 / 67.30 52.82 / 63.42
Ours (PV+YOLO+GDino) 67.28 / 73.97 60.36 / 68.01 54.31 / 63.58
TABLE III: Perfomance on the Average Precision (AP) metric for our method on all difficulty sets on KITTI-val for the Pedestrian class. Our late fusion method consistently outperforms the original specialized methods in 3DAP across all the difficulty levels. The color coding in yellow emphasizes the degree of improvement on 3DAP over the baselines.
@@ V-B Average Precision
Late fusion with general and open-vocabulary camera models (YOLO-NAS and GroundingDINO) perform as well as the tested special-purpose camera model (MonoDETR). This is surprising since the baseline 2D-AP of these detectors is much lower than the specially trained model. This provides evidence that general-purpose camera detectors can be used to enhance special-purpose LiDAR detectors.